### v2

# Neural timescales from a computational perspective

Abstract
Neural activity fluctuates over a wide range of timescales within and across brain areas. Experimental observations suggest that diverse neural timescales reflect information in dynamic environments. However, how timescales are defined and measured from brain recordings vary across the literature. Moreover, these observations do not specify the mechanisms underlying timescale variations, nor whether specific timescales are necessary for neural computation and brain function. Here, we synthesize three directions where computational approaches can distill the broad set of empirical observations into quantitative and testable theories: We review (i) how different data analysis methods quantify timescales across distinct behavioral states and recording modalities, (ii) how biophysical models provide mechanistic explanations for the emergence of diverse timescales, and (iii) how task-performing networks and machine learning models uncover the functional relevance of neural timescales. This integrative computational perspective thus complements experimental investigations, providing a holistic view on how neural timescales reflect the relationship between brain structure, dynamics, and behavior.
1 Introduction
Ongoing neural activity unfolds across a wide range of timescales. These timescales are often characterized by the decay rate of a neural signal’s autocorrelation function. While precise definitions differ, recent experimental studies of neural timescales present a consistent picture across different recording modalities, species, and cognitive tasks: timescales increase along the cortical hierarchy, are diverse within each brain area, and are often correlated with behavioral variables during cognitive tasks (for recent reviews of empirical evidence, see [1, 2, 3]). As such, this fundamental quantity—the characteristic time constant of neuronal dynamics—may be an important signature of neural circuits across the brain. Importantly, neural timescales have been suggested to take on a “representational” role, i.e., the diversity of timescales across the brain mirrors the multiple fluctuation timescales of relevant information in a dynamic environment, which is leveraged by the brain for perception, memory, planning, and action [1, 2, 3]. However, while a large body of experimental observations demonstrates the relevance of neural timescales, crucial differences exist between how these timescales are defined and measured. Furthermore, it is experimentally challenging to probe the mechanisms underlying how neural timescales arise, and whether they are necessary (or sufficient) for representation and transformation of information in the brain.
Supplementing experimental investigations, we here contribute a complementary perspective and review how computational methods have shed light on the origins and functions of neural timescales, while clarifying definitions for different types of timescales in the literature (Box 1). Specifically, we highlight three directions for how computational methods can disentangle different types of timescales and distill the broad set of empirical observations into quantitative and testable theories, formulated as the following questions (Fig. 1):
  • “Neural timescales” are measured from different brain signals, during different tasks, and quantified using different methods: how do these measurements relate and differ from one another? (Section 2)
  • Which cellular and circuit processes contribute to shaping measured timescales, and how can mechanistic models of brain dynamics help propose and exclude candidate mechanisms? (Section 3)
  • When are “representational neural timescales” that mirror environmental timescales necessary (or sufficient) for supporting task-relevant computations, and how to use functional models from machine learning to explore such hypotheses? (Section 4)
Moreover, we review how recent advances in machine learning (ML) and artificial intelligence (AI) have opened new avenues to develop novel data analysis methods and computational models, getting us closer to an understanding of how neural timescales capture the relationship between brain structure, dynamics, and behavior.
[!ht] Box 1. Glossary: Different usages of “timescales” The definition of timescale varies in the literature across studies. Here, we briefly outline different notions of timescale that are commonly (and sometimes implicitly) used, with more detailed discussions throughout the Review:
Biophysical time constants:
Timescales can be thought to correspond to the passive membrane or adaptation time constants of single neurons. Similarly, timescales measured from field potential data may relate to time constants of synaptic or ion channel currents (Box 2).
Activity timescales in linear(ized) networks:
Another common notion refers to the decay time constant of population activity. Specifically, the output of a linear time-invariant system (i.e., a linear recurrent neural network) can be decomposed into independent modes of (complex) exponentials with different decay time constants (Box 3). This notion is often leveraged to interpret estimated timescales with the implicit assumption that network dynamics is (locally) linear, since “time constants” in nonlinear networks are state-dependent and thus ill-defined as a global, fixed property.
Response timescales:
Aside from the above “intrinsic” properties of the neural system under observation, measured timescales can also reflect timescales embedded in the input or be modulated by task and behavioral states. Response timescales could thus correspond to multiple timescales in a streaming sensory stimulus (e.g., a movie), or more generally, to upstream neural activity onto the downstream, recorded region (e.g., V1 to V4). They can also arise from modulations by top-down inputs during a behavioral task.
Any quantification of timescales from neural recordings is likely to reflect, to different degrees, a mixture of the above. Therefore, a key challenge moving forward for both computational methods and experimental design lies in separating those properties of neural systems from measurements of neural activity (Section 2), as well as understanding their roles in neural circuit dynamics and computation (Sections 3 and 4).
Fig. 1: Graphical summary. This review discusses three main directions, demonstrating conceptually: (i) how estimation methods contribute to advancing the measurement of timescales from different experimental paradigms (e.g., recording modalities), (ii) how mechanistic models can help in uncovering the underlying circuit origins of different timescales, and (iii) how functional modeling facilitates investigation of their computational relevance.
2 Estimating timescales from neural recordings
Neural timescales have been estimated from data collected across various experimental paradigms, including different species, tasks, and recording modalities, as well as using different estimation methods (Table 1). While studies of neural timescales often concentrate on the similarity of their findings across different paradigms and methods, aforementioned differences also introduce assumptions and biases on estimated timescales that may lead to important conceptual deviations. In the following, we discuss these assumptions and biases and how they should be incorporated when assessing the generalizability of results.
Fig. 2: Diversity of neural timescales across and within brain areas. a. Average estimated timescales from spontaneous neural activity measured by different recording modalities consistently increase along the cortical hierarchy, but their values differ by orders of magnitude (data from [4, 5, 6]). b. Neural timescales are also diverse across individual neurons within one brain area (data from [7]). MT: medial-temporal area, LIP: lateral intraparietal area, OFC: orbitofrontal cortex, LPFC: lateral prefrontal cortex, DLPFC: dorsolateral prefrontal cortex, ACC: anterior cingulate cortex.
2.1 The impact of experimental paradigm on estimated timescales
One of the most robust observations across different studies is the hierarchical organization of neural timescales, i.e., timescales being faster in primary sensory and motor areas, and slower in transmodal association areas. The hierarchy has been reported across various species (rodents and primates), recording modalities (spiking activity, calcium imaging, LFP, ECoG, and fMRI), and cognitive tasks (Table 1), suggesting that shared principles may give rise to these observations. However, there are also important distinctions that may be useful for uncovering the underlying mechanisms.
Modality Method Model-free Variants of exponential fit Autoregressive models Bayesian methods
Spiking activity Mendoza-Halliday et al. [8],  Wang et al. [9],  Meirhaeghe et al. [10],  Tang et al. [11] Zeisler et al. [12]∗∗,  Murray et al. [4],  Wasmuht et al. [13],  Cavanagh et al. [7],  Fascianelli et al. [14], Cirillo et al. [15],  Bernacchia et al. [16],  Fontanier et al. [17]∗∗,  Zeisler et al. [12]∗∗,  Siegle et al. [18]∗∗,  Rudelt et al. [19]∗∗,  Rudelt et al. [20],  Zeisler et al. [12]∗∗ Spitmaan et al. [21]∗∗,  Song et al. [22]∗∗,  Song et al. [22]∗∗ Zeraati et al. [23],  Zeraati et al. [24]∗∗,  Neophytou et al. [25],  Gonçalves et al. [26],  Gao et al. [27],  Vinogradov et al. [28]
Calcium imaging Çatal et al. [29]∗∗ Runyan et al. [30],  Imani et al. [31],  Piasini et al. [32]∗∗ Pinto et al. [33]∗∗ Nair et al. [34], Vinograd et al. [35],  Vinogradov et al. [28]
LFP, ECoG, EEG Honey et al. [36] Donoghue et al. [37],  Brake et al. [38], Chaudhuri et al. [39],  Gao et al. [5]∗∗,  Çatal et al. [29]∗∗, Cusinato et al. [40],  Gao et al. [5],  Manea et al. [41]∗∗
fMRI Hasson et al. [42],  Raut et al. [43],  Fallon et al. [44],  Watanabe et al. [45],  Shafiei et al. [46], Manea et al. [6],  Sethi et al. [47] Fallon et al. [44]
Table 1: Overview of timescale measurements in experimental literature. Timescales are measured using various recording modalities, estimation methods, and species (indicated by colors as human, non-human primate and rodent). * indicates studies that measured timescales from task-driven activity and ** from both spontaneous and task-driven activity. The rest measured timescales from spontaneous activity. Repeated entries indicate multiple methods, modalities, or species.
Timescales estimated from different recording modalities increase along the cortical hierarchy, though their absolute values differ by orders of magnitude (Fig. 2a). These differences may stem from either biophysical properties of the circuits (e.g., time constants), or constraints in the experimental recording setup, such as sampling rate. Spiking activity timescales are often reported to be in the range of tens to hundreds of milliseconds [4, 21], whereas fMRI BOLD timescales are in the range of a few seconds [43, 44, 6]. Since BOLD signals partially reflect population spiking activity [48, 49] by measuring a lagged hemodynamic response [48], timescales estimated from fMRI may correspond to “slowed” population spiking timescales. Estimated timescales from LFP [41] and ECoG [5], on the other hand, are approximately ten times faster than spiking activity in similar brain regions, which seems to be at odds with the notion that LFP and ECoG are slower aggregate signals [48, 50]. This observation suggests that timescales estimated from different modalities may be dominated by distinct but potentially correlated physiological processes, e.g., fast transmembrane and synaptic currents versus slow recurrent network activity. In fact, when explicitly accounting for multiple timescales in the estimation method [23], both fast and slow timescales can be uncovered from spiking activity [24]. At the same time, currently used slow fMRI scan rates may further limit the detection of faster timescales in neural activity [51, 52]. Similar issues can affect timescales measured from calcium imaging: The time constant of calcium release, as well as the time constant of the calcium indicator and the imaging scan rate, can cause the same underlying neural dynamics to appear as having a different timescale [53]. Thus, when interpreting their potential functional relevance, it is important to distinguish whether estimated timescale differences arise from underlying neurophysiological properties, measurement process constraints, or both.
Neural timescales can also be estimated from different behavioral states and task conditions, resulting in two conceptually different quantities: (i) intrinsic timescales reflecting spontaneous fluctuations of neural activity (e.g., during fixation or rest), and (ii) response timescales arising from stimulus- or task-driven activity fluctuations. While both intrinsic and response timescales increase along the cortical hierarchy [4, 18, 20, 32, 6, 5, 41, 43, 21, 22, 54, 36, 42], they are not correlated within single neurons [21]. The reason behind this observation may be two-fold: First, single-neuron activity timescales are highly diverse within individual brain areas [7, 13, 14, 15], and even areas at the top of the hierarchy have many neurons with fast timescales (Fig. 2b). Various studies have shown that neurons with fast and slow timescales engage in different task epochs and are relevant to distinct types of computation (e.g., [7, 13, 14, 17], for a review see [1]). Second, neural timescales are state- and input-dependent, and change according to task demands [5, 24, 41, 17] or input timescales [32]. Thus, distinct mechanisms may underlie intrinsic and response timescales. In the following sections, we discuss how generative model-based methods can incorporate modality- and task-specific components when estimating neural timescales and how mechanistic models allow us to disentangle their underlying mechanisms, e.g., state-dependent timescales arising from nonlinear brain dynamics.
2.2 Different methods for estimating timescales
Extracting timescales from neural activity usually requires two steps: (i) computing a form of summary statistic that quantifies temporal fluctuations, and (ii) estimating timescales from the summary statistic (Fig. 3). For the summary statistic, most methods use some variant of time-lagged metrics, such as autocorrelation (AC, or its frequency domain equivalent, power spectral density, PSD) or delayed mutual information (MI, Fig. 3a, top). For example, intrinsic timescales are characterized by the decay rate of the spontaneous activity autocorrelation computed from individual experimental trials [4, 7, 13, 14, 16, 17, 18, 19, 24, 23, 5, 31, 32, 37, 39, 29, 6, 41, 44, 43, 45], whereas response timescales are quantified by the autocorrelation of trial-averaged activity post-stimulus [32, 18]. Response timescales can also be estimated with decoding methods that assess how long the encoding of a given stimulus or task event persists over time [30, 20, 19, 32]. Here, the summary statistic is often constructed using metrics such as decoding accuracy or predictability as a function of time (Fig. 3a, bottom) [30, 20, 19].
Fig. 3: Different methods for estimating neural timescales. a. Neural activity fluctuations can be characterized by various summary statistics such as the autocorrelation or delayed mutual information (AC / MI, top), or information content in neural activity about a preceding stimulus (bottom) (adapted from [30, 20, 19]). b. Methods for estimating timescales ($\tau$) from summary statistics generally vary along two axes: (i) Computational complexity (vertical), (ii) Number of features they capture in data (horizontal): from the approximate decay rate to capturing autocorrelation shape by including a baseline representing slow processes across trials or oscillatory components, to including task-related components and prior knowledge about the data (panels are adapted from [45, 4, 55, 21, 23]).
Using a computed summary statistic (e.g., AC or MI) as input, methods for estimating timescales rely on distinct theoretical assumptions and incorporate different features of neural dynamics. Here, we catalog existing methods along two axes: (i) computational complexity (often as a result of using more detailed models) and (ii) the number of features they capture in data (Fig. 3b). Note that, while in the following we mainly focus on quantifying timescales from the autocorrelation function, similar analyses can be applied to information-theoretic summary statistics.
The first group of methods does not assume any strict mathematical form for the autocorrelation shape. These methods characterize the decay rate as, for example, the time lag where autocorrelation drops to $1/e$ [56] or 0.5 (i.e., half width at half maximum [43]), or the integral of positive autocorrelation values [45, 6]. We thus refer to these methods as “model-free”. Computing such measures is straightforward, and they are typically estimated directly from the data. However, such model-free methods do not consider the full shape of data autocorrelation. As a result, in the presence of complex temporal features in neural activity beyond a single exponential decay, different methods can give different results on the same data (for example, see [56, 57]).
The second group of methods assumes a mathematical function for the autocorrelation to capture the relevant data features (i.e., model-based), and estimate timescales from the parameters of the fitted function (e.g., via least squares). The simplest assumption is that the autocorrelation of neural activity follows an exponential decay function, $\textrm{AC}(t)=e^{-t/\tau}$, where decay time constant $\tau$ reflects the timescale of underlying neural dynamics. This is equivalent to the assumption of obtaining a Lorentzian function as a power spectral density in the frequency domain, i.e., $\textrm{PSD}(f)\sim(f^{2}+f_{k}^{2})^{-1}$, where the knee frequency $f_{k}$ is translated to timescale as $\tau=(2\pi f_{k})^{-1}$ [5]. The exponential decay (or Lorentzian function) can be further augmented with additional features to better capture the complexity of temporal dynamics. For example, the autocorrelation of neural dynamics can exhibit multiple timescales [39, 21, 24] that may correspond to distinct neural mechanisms. In such cases, timescales can be estimated by fitting a linear mixture of exponential (or Lorentzian) functions [39, 19]. Moreover, very slow timescales can also arise from firing rate fluctuations across different trials, which can be accounted for by fitting the autocorrelation with an exponential decay function and an additional offset term corresponding to longer-term processes (e.g., [4, 7], see [58] for a toolbox). Neural activity can additionally contain oscillatory dynamics that appear as periodic fluctuations in the autocorrelation or peaks in the PSD, which can impact the estimated timescale or knee frequency, respectively. To overcome this problem, methods such as Spectral Parameterization [37, 5, 38] model oscillatory peaks as a mixture of Gaussians, which are removed before fitting an extended form of the Lorentzian function as $\textrm{PSD}(f)=C(f^{\chi}+f_{k}^{\chi})^{-1}$ (i.e., with a variable 1/f power-law exponent) to allow more flexible fitting of the neural power spectra.
Finally, a more accurate estimation of timescales can be obtained using a generative process capturing the statistics of neural time series [23, 21, 25, 59]. Such methods are particularly powerful since they can explicitly incorporate modality-specific properties and task variables. For instance, including task variables in an autoregressive generative process has allowed for disentangling intrinsic and response timescales such as fluctuations induced by reward [21, 33, 22]. Moreover, multiple timescales can be explicitly modeled in the generative process, for instance, as a mixture of autoregressive processes with distinct timescales [23, 24, 21]. Generative models can also capture experimental limitations, such as short trial durations that affect the shape of data autocorrelation and can lead to systematic underestimation of timescales by model-free and model-based methods discussed earlier [23].
The parameters of the generative process can be directly fitted to neural time series [21] (e.g., via least squares or maximum likelihood) or can be used to instantiate a generative model in Bayesian inference methods [23, 59, 25, 26, 60, 61, 27, 62]. Bayesian methods aim to uncover the whole set of parameters that is compatible with the data (i.e., the posterior distribution) instead of obtaining the single best-fitting value of timescales (or other parameters), allowing them to better deal with the stochasticity of neural activity. For simple models, the posterior distribution can be computed analytically [59], but for more complex models, simulation-based methods such as approximate Bayesian computation (e.g., abcTau [23]) and simulation-based inference (SBI [26, 60, 61, 63, 27]) can be applied to obtain approximate posterior distributions. In such methods, synthetic time series are produced from an explicit (statistical or mechanistic) generative model mimicking similar statistics as neural recordings. Then, by matching the summary statistics (e.g., autocorrelation, PSD, or time series) of synthetic data to the experimental data, we can estimate timescales using the parameters of the matched generative model.
The generative model and summary statistics can be chosen flexibly, allowing them to capture different data features. It is possible to fit different generative models (e.g., models with different numbers of timescales or oscillatory components) and use Bayesian model comparison techniques to find the best models describing the neural data [23, 64]. Since the generative model can also incorporate modality-specific properties, such as timescale ranges, stochastic elements (e.g., spiking activity as a point process versus LFP as a continuous signal [65, 66, 67]), or even particular biophysical properties, it can be used to uncover possible models describing each modality. Therefore, while these methods often come with additional computational costs from performing simulations and posterior inference, they can incorporate important mechanistic assumptions and data features—both based on neurobiology and biophysics of the signal-generating process—when estimating timescales.
3 Mechanistic models of timescales in the brain
Mechanistic models, when constrained by experimental observations of neural timescales, can provide insights into how the diversity of timescales arises within and across different brain areas. Such models have so far been used to explain the following experimentally observed properties of neural timescales: (i) the hierarchical structure of neural timescales across different brain areas [68], (ii) the diversity of timescales within one brain area [69, 70, 24], and (iii) the flexibility of timescales to change according to different task variables [24, 9, 71, 72]. Mechanisms for capturing diverse intrinsic timescales are generally categorized into cellular and synaptic mechanisms on the level of individual neurons and network-mediated mechanisms through interactions between neurons (Fig. 4). At the same time, other processes, such as bottom-up (e.g., sensory) or top-down (e.g., from higher cortical areas) inputs and neuromodulatory mechanisms, can dynamically modulate timescales and give rise to stimulus- and task-dependent response timescales. Moreover, different models capture neural timescales on different spatial scales of observations, e.g., timescales of single neurons versus population activity. Thus, comparing these models can reveal how neural timescales across different scales and modalities can be related to each other.
3.1 Cellular and synaptic mechanisms
In computational models, timescales are often controlled by incorporating various biophysical time constants of cellular and synaptic processes in the dynamics of individual neurons that reflect heterogeneous properties of ionic channels and synapses [73, 74]. Most dynamical system models include a time constant in the dynamical equation of each neuron, interpreted as the membrane time constant ($5-50$ ms [75]) (Box 2). Slower timescales can be implemented by introducing other biophysical features such as slow synaptic time constants [76, 56, 77], timescales of short-term synaptic depression and facilitation [78, 79], or adaptation mechanisms [80, 81, 82, 56]. Spike-frequency adaptation, in particular, spans a wide range of fast and slow timescales [83], which could follow a power law [82]. Diverse adaptation timescales are suggested to lead to efficient coding in individual cortical neurons driven by behaviorally relevant signals [82] and facilitate computations on temporally dispersed inputs [84].
Box 2. Activity timescale as an estimated quantity vs. time constant as a parameter “Neural timescales” in the literature, for the most part, refer to the estimated autocorrelation decay timescale of measured neural activity, i.e., by describing the autocorrelation function of spike counts or LFP time series as $\textrm{ACF}(t)=e^{\frac{-t}{\tau_{\textrm{fit}}}}$. Methods discussed in Section 2 generally aim to find the value of $\tau_{\textrm{fit}}$ (or multiple $\tau_{\textrm{fit}}$s) that best fit the data. On the other hand, mechanistic (i.e. Hodgkin-Huxley) and functional (i.e., rate-based recurrent neural networks) models in Sections 3 and 4 typically include a time constant parameter that describes how the impact of input into a neuron decays over time, e.g., $\tau_{\textrm{param}}\frac{dV}{dt}=-(V(t)-V_{\textrm{rest}})+I(t)$, where $I(t)$ includes time-varying bottom-up (stimulus-driven), top-down, and recurrent inputs. It is important to note that the former, estimated $\tau_{\textrm{fit}}$ does not necessarily need to correspond to the latter, underlying model parameter $\tau_{\textrm{param}}$, and in general, does not. Section 3 outlines multiple mechanisms, including those unrelated to time constants, such as network connectivity, that contribute to shaping $\tau_{\textrm{fit}}$ we observe in data (see also Box 3).
Fig. 4: Mechanisms underlying generation and modulation of neural timescales. Computational models show that neural timescales are shaped by cellular biophysical properties (left, e.g., membrane, synaptic, and adaptation time constants) and network interactions (right, e.g., spatially structured or clustered connectivity, inhibition). Additionally, timescales can be modulated (bottom) by neuromodulatory signals and sensory inputs from the environment. SFA: spike-frequency adaptation; STD/F: short-term synaptic depression/facilitation.
However, biophysical time constants at the single-neuron level cannot completely explain the temporal structure of neural dynamics. For instance, a Hodgkin-Huxley model of an isolated neuron with detailed dynamics of membrane potential and ionic and synaptic currents cannot capture the full autocorrelation shape of neural activity recorded in prefrontal areas [17]. Biophysical time constants alone define the autocorrelation timescale of neural activity—henceforth referred to as activity timescale (Box 2)—when a neuron is isolated, but when neurons are embedded in a network, their activity timescale may differ significantly from their biophysical time constants. At the same time, distinct biophysical mechanisms may interact differently in the presence of network interactions compared to isolated neurons. In particular, while activity timescales of network-embedded neurons increase with the synaptic time constant [56], their relationship to the adaptation time constant is less straightforward: Analyzing similar models of randomly connected excitatory and inhibitory units, Beiran and Ostojic [56] reported that activity autocorrelation timescales do not depend on the adaptation time constant, whereas Muscinelli et al. [57] report a positive relationship. It is important to note, however, that the autocorrelation of neural activity in networks with adaptation does not follow a simple exponential decay. One possible explanation of the apparent disagreement in results could lie in the different methods used to measure timescales in each study, thus highlighting the need to understand how timescale estimation methods impact the conclusions we draw from empirical data and computational simulations. In the next section, we review how incorporating network mechanisms, in addition to cellular and synaptic mechanisms, could help to better explain observations of diverse and longer neural timescales.
3.2 Network and connectivity mechanisms
Neurons in the brain are embedded within rich networks [85, 86] that affect the correlation patterns of neural dynamics [87, 88, 89, 90]. Therefore, a mechanistic understanding of neural timescales requires the context of interactions within a network. Most models rely on recurrent (excitatory or inhibitory) connectivity to explain how diverse neural timescales emerge, but as we discuss later, feedforward networks with rich temporal inputs can also give rise to diverse and hierarchical timescales [91].
Recurrent connectivity (within or across brain areas) can create positive or negative feedback that controls activity timescales. This can be done, for example, through creating attractor dynamics that give rise to persistent activity (e.g., line attractors) [92, 93] or shifting the network dynamical state closer to a critical transition [94, 95, 68, 39, 24, 96, 97, 98, 99, 100]. By analyzing linear recurrent networks (i.e., linear dynamical systems), we can define and compute the activity timescales of neurons analytically based on eigenvalues of the connectivity matrix (Box 3). However, in nonlinear networks, such timescales are not uniquely defined and can only be specified for a given (linearized) dynamical state, such as steady-state dynamics around fixed points [93, 101, 102, 34, 35].
The strength and structure of recurrent excitation influence the amount of positive feedback in a network and, consequently, the autocorrelation timescale of neural activity. In general, timescales increase with the strength of excitation [103, 68, 104, 105, 106, 94, 95, 68, 39, 24]. Moreover, heterogeneity in connection strengths can give rise to distinct timescales across network units [69, 70, 107, 100, 16]. For fixed connection strengths, timescales can additionally be affected by the connectivity structure. For instance, clustered connectivity creates metastable dynamics, and noise-induced transitions between these states give rise to slow timescales [108, 109, 110]. Metastable dynamics can also emerge in homeostatically regulated spiking networks [111].
Box 3. Activity timescales of a linear network As an illustrative example, we review how to compute the activity timescale of a linear recurrent network (i.e., linear dynamical system). Consider the system $\frac{dx(t)}{dt}=Ax(t)$ where $x(t)$ is the state vector, e.g., neural activity at time $t$, and $A$ is a square connectivity matrix. To acquire the timescales of the system dynamics, we solve for the eigenvalues of the matrix $A$, $\lambda_{i}$, through the characteristic equation, $A\textbf{u}_{i}=\lambda_{i}\textbf{u}_{i}$. In general, eigenvalues are complex and have the form $\lambda_{i}=\textrm{Re}(\lambda_{i})+j\textrm{Im}(\lambda_{i})$. For a stable decaying mode, i.e., non-exploding, the real part of the associated eigenvalue must be negative, and its decay timescale is defined as $\tau_{i}=-\frac{1}{\textrm{Re}(\lambda_{i})}.$ The dominant activity timescale of the entire system (i.e., estimated as $\tau_{\textrm{fit}}$) is then typically determined by the mode whose eigenvalue has the largest real part. Furthermore, for a linear network composed of units with fixed membrane time constant $\tau_{\textrm{param}}$, the system dynamics follow $\tau_{param}\frac{dx(t)}{dt}=Ax(t).$ Assuming $\tau_{\textrm{param}}$ to be the same for all neurons, the activity timescale is scaled accordingly, $\tau_{i}=-\frac{\tau_{\textrm{param}}}{\textrm{Re}(\lambda_{i})}.$ This approach has been applied to connectivity matrices incorporating further biological constraints, e.g., random networks obeying Dale’s law [219]. For nonlinear networks, it is necessary to first linearize the system around a specific point in state space before conducting the above analysis (see, e.g., [102]).
Furthermore, spatially structured connectivity leads to multiple dominant timescales in local dynamics, creating autocorrelation shapes beyond a single exponential decay [112, 24]. In these networks, each timescale is created by neural interactions on different spatial scales, such that slower timescales arise from interactions on a larger spatial scale, consistent with similar observations and theories on the spatiotemporal scale of oscillatory brain activity [113].
Negative feedback via recurrent inhibition also plays an important part in forming neural timescales. Strong inhibitory-to-inhibitory connections allow for slow neural timescales required for working memory maintenance in a recurrent network model [114]. Moreover, in a detailed biophysical model of excitatory and inhibitory neurons, inhibition controls the stability and transitions of metastable states, giving rise to slow timescales [17]. In this model, the full shape of neural autocorrelations in frontal areas is explained by a combination of cellular (i.e., after-hyperpolarization potassium and inhibitory GABA-B conductance) and network mechanisms. In addition, recurrent inhibition together with membrane and synaptic time constants can account for distinct temporal dynamics in spiking activity and LFP [115].
3.3 Input-driven and modulatory mechanisms for dynamically shaping timescales
The mechanisms discussed so far explain how a specific range of timescales can emerge due to “static” properties of single neurons and networks. However, neural timescales are themselves dynamic, and can adapt to task demands [5, 23, 17, 9, 41, 72, 32, 71]. Here, we discuss how external inputs, including bottom-up sensory inputs and top-down inputs from higher brain areas, as well as neuromodulators such as acetylcholine (ACh) can alter neural timescales.
External input can have both direct and modulatory effects on the resulting timescales of a population. At the cellular level, incoming input can alter biophysical properties such as by bringing individual neurons to a high-conductance state (lower resistance), which reduces the membrane time constant [116] and facilitates the distinction of distant synaptic inputs [117]. The spatial and temporal statistics of the input itself can also impact response timescales measured from neural activity. Different features of sensory inputs can vary at different timescales (e.g., moving edges vs. persistent object identity). For example, theoretical work using feedforward models constructed based on slow feature analysis has suggested that fast input timescales are encoded in lower levels of visual hierarchy, while slower timescales are encoded in higher levels [91, 118, 119, 120, 121] (for a review, see [122]). In these models, learning invariant or slowly varying features of sensory stimuli gives rise to progressively slower representations and temporal stability of the stimulus. These model predictions have been confirmed experimentally in the rodent visual hierarchy [32]. Similar behavior can arise in recurrent networks, where populations with fast (slow) intrinsic timescales respond stronger to faster (slower) components of the input [70]. Furthermore, the spatial scale of stimuli modulates the timescale of population spiking activity by affecting feedforward and recurrent interactions [123]. Thus, depending on the spatiotemporal profile of the input, different timescales thereof may be encoded and reflected in neural activity.
Even without embedded temporal dynamics, input can modulate activity timescales by shifting the dynamical state of nonlinear networks. As discussed above, timescales can be estimated for nonlinear networks at a specific state. Different input levels thus place the network in different states, which can alter the “local” timescales [112, 93]. This mechanism could explain the modulation of neural timescales during selective-attention [24], aggression [34], and timing tasks [9, 71]. Moreover, it can capture how the organization of brain-wide response timescales during tasks deviate from the expected hierarchy of intrinsic timescales [93]. An interesting example is the sharp transition in response timescales from the middle temporal (MT) area to the medial superior temporal (MST) during working memory delay, despite their closeness in the visual hierarchy [8]. Large-scale cortical models suggest that this transition arises from nonlinear interactions between different brain areas, where distinct intrinsic and delay-activity response timescales are shaped by dynamics in different fixed points [93].
Neuromodulatory projections from diffuse subcortical nuclei can also alter timescales as a result of their impact on cellular and network properties [124, 125, 126]. ACh, for instance, can change synaptic time constants and lead to persistent firing (e.g., during pain processing [127]). An increase in ACh, such as during selective attention, can also strengthen the thalamocortical synaptic efficacy by affecting nicotinic receptors, while weakening lateral cortical interactions by affecting muscarinic receptors [128]. Similarly, the different contribution of glutamatergic receptors in feedforward and recurrent pathways affects the time course of neural population responses in distinct ways [129]. Thus, neuromodulators can indirectly adapt neural timescales to task demands through selective gain modulation.
Altogether, a combination of cellular, network, input, and neuromodulatory mechanisms can shape the diversity and flexibility of neural timescales. We argue that constraining mechanistic computational models with realistic timescales can provide new insights into circuit mechanisms underlying neural population dynamics. While such models are typically challenging to constrain, recent open-source toolboxes leveraging approximate Bayesian computation (ABC) [23] and simulation-based inference (SBI) [60, 61] have enabled efficient discovery of data-consistent mechanistic models [26, 23, 130, 27, 28, 131]. Based on the available experimental measurements, one can constrain models with a single estimated timescale (e.g., average timescale of a brain area in Fig. 2a to average timescale of a network model), or with a distribution of timescales (e.g., diverse timescales within an area Fig. 2b to distribution of timescales within a model). As we discuss next, assessing for realistic timescales (e.g., similar to realistic firing rates) also allows us to test the plausibility of different models aimed to explain computation and function.
4 Timescale as signatures of dynamics supporting network computation
Finally, we return to function: Recent experimental studies have consistently observed that timescales of both single-unit and population dynamics correlate with measurements such as behavioral timescales and task performance [1, 2, 3]. Neural timescales are thus implicated to be a signature of dynamical properties necessary for relevant computations. However, to make causal claims about the necessity and sufficiency of dynamics operating on different timescales for network computation, perturbation experiments are required to assess the functional impact of timescales over and above, e.g., firing rate modulations, which can be challenging in vivo.
Recently, task-performing artificial neural network (ANN) models—including spiking and rate-based recurrent neural networks (SNNs and RNNs), as well as deep neural networks in machine learning—offer a viable, complementary direction: Such functional models circumvent the practical challenges of manipulating timescale-related variables in the in vivo brain, while enabling us to probe loss- or gain-of-function in experimental task paradigms. In computational and cognitive neuroscience, these networks are typically designed to mimic the input-output behavior observed in animal or human participants performing a task [132, 133] (Fig. 5a). To match the target output, the ANN parameters are either “hand-crafted” or analytically chosen when possible (e.g., for simple tasks; Fig. 5b, left), or “trained” using an optimization algorithm minimizing a loss function (e.g., for larger networks and more complex tasks; Fig. 5b, middle). Tuned networks can then be further analyzed and perturbed to generate hypotheses for the mechanisms of computation at hand, in particular with respect to properties of interest in phase space (Fig. 5c) [134, 135] and timescales of population dynamics (Fig. 5d).
In this section, we review two broad categories of timescale-related investigations using ANNs: First, we discuss those that study neural timescales as an emergent characteristic of task-performing networks, including both classical works on dynamical properties of “hand-tuned” networks and recent ones employing task-optimized networks. Second, we discuss recent studies that explicitly optimize time constants and timescale-related parameters to enhance computational capacity in artificial networks (Fig. 5b, right), including biologically plausible SNNs and RNNs, as well as more powerful models in deep learning applications.
Fig. 5: Task-performing models reveal functional relevance of timescales. a. Network models are set up to achieve input-output transformations similar to tasks performed by animals and humans. b. Model parameters, such as network weights and neuronal time constants, can be manually tuned when feasible or computationally optimized via gradient descent. c. A certain geometry of population dynamics is desirable for network computation, such as fixed points, line attractors, and limit cycles. d. Activity timescales and fitted time constants in network models match experimental data.
4.1 Emergent timescales in task-performing neural networks
Computational neuroscience has had a long-standing interest in the repertoire of dynamics—and by extension, their timescales—to elucidate how neural circuits can represent and perform tasks with complex temporal structures. For instance, in classical hand-tuned networks, fixed points, line attractors, and limit cycles (Fig. 5c) are dynamical motifs that can be leveraged for memory [136, 137], input integration and decision-making [138, 139, 140], and timing [141], respectively (see [142] for a recent review on attractor models of neural computation). Beyond the geometry itself, timescales of the latent dynamical system near and on these low-dimensional manifolds are also key properties of interest. Crucially, those timescales govern, for example, how quickly and how stably a network can integrate new input while remembering or forgetting old ones, filter out fast, noisy fluctuations to sense slower signals, and receive inputs of different frequencies. More recently, and consistent with hand-tuned networks [143], task- and activity-optimized models also exhibit fixed points corresponding to persistent inputs or memories [144], line attractors for tracking slow environmental or behavioral variables [34, 35, 145, 146], and (quasi-)limit cycles traversed at different speeds corresponding to the timing of network inputs or outputs [9, 71].
Additionally, related research on reservoir computing, including liquid state machines [147] and echo state networks [148], similarly considered the activity timescales of RNNs from the perspective of latent dynamical systems, since they require long-term but not divergent (or exploding) dynamics at the edge-of-chaos [149, 150]. While reservoir networks typically have trained input and output weights, the recurrent weights are fixed due to the challenges of propagating errors over many time steps. As such, much attention was dedicated to investigating how weights can be thoughtfully initialized to maximize the network’s computational capacity [151, 152], and in particular, how input-driven chaos in such networks can be tuned to represent a flexible repertoire of timescales (or temporal filters) for time-dependent computations [149].
Notably, an appropriate distribution of timescales is not only important for implementing these computations, they are useful for identifying key dynamical structures in the first place. For instance, an asymmetry in the decay time constants of a (locally) linear system (Box 3) is indicative of one slow direction and many fast directions—in other words, a line attractor [153, 34]. Identification of such behaviorally-relevant latent structures in network models further enables causal manipulation of the neuronal populations involved, closing the loop between theory and experiment in dissecting circuit mechanisms underlying complex behaviors such as the role of subcortical brain structures in mating and encoding of affect in self and others [35, 145, 146].
Alternatively, one could directly compare the activity timescales of individual neurons in experimental recordings with that of neurons in task-performing SNNs and RNNs (Fig. 5d). This approach can similarly aid in understanding cellular, network, and learning-related mechanisms of computations underlying e.g., temporal scaling [9], working memory [114, 154], and other motor or cognitive tasks [155]. These ideas can be further extended to study networks endowed with more sophisticated and biologically-realistic constraints such as Dale’s law [156, 114] and dendritic compartmentalization [157], hypothesized computational mechanisms such as oscillatory phase coding [158, 159], as well as better analytical tractability (e.g., low-rank RNNs [160]), improved expressiveness, and trainability [161, 162, 163]. Targeting the same sought-after dynamical repertoires, “hybrid” approaches can leverage more sophisticated weight initialization or regularization strategies to nudge the network toward line attractors or limit cycles after training [164, 163]. Furthermore, since population activity timescales reflect the state of neural dynamics relative to critical transitions, timescales have also been used to study how the state of dynamics in spiking neuromorphic networks affect the performance of tasks with varying complexity [165].
Altogether, these studies provide insights into specific tasks or network architectures underlying computations of interest from the perspective of cellular and network dynamics. Leaning on the notion that the geometry of representation in a neural population is intimately linked to (the timescales of) its dynamics, and that the latter may be even more important than the former [166], both classical and recent works lend support to the importance of diverse neural timescales for different computations. In other words, the temporal signatures of a dynamical system are a reflection of its phase portrait geometry, and constraining one to satisfy certain computational requirements often means constraining the other as well [162].
Nevertheless, it may be possible for networks to acquire similar low-dimensional geometries but vastly different timescales of dynamics. Therefore, we suggest explicit comparisons of in silico and in vivo timescales in task-performing networks as a straightforward yet synergistic research direction to provide explicit evidence for the necessity (or lack thereof) of different timescales of dynamics. It should be noted, however, that similar to nonlinear biophysical networks discussed in Section 3.2 and Box 3, analyses of timescales in the context of task-performing models similarly rest on the assumption of linearized systems in a particular state. While such analyses provide insights on “local” dynamics and computational properties, individual timescales do not inform of global properties of the network, which may require sampling a distribution of timescales across multiple states and task contexts.
4.2 Optimization of timescale-related parameters
Another powerful way to use modern ANNs to study the effect of neural timescales on computation is to directly optimize timescale-related parameters in the network model, such as synaptic, membrane, and adaptation time constants, or more phenomenological variables like temporal receptive field size and context length:
In a standard RNN, all neurons are given a single fixed time constant that is roughly analogous to the membrane time constant of biological neurons, which determines how quickly the effect of an input on that neuron decays. In reality, neurons across the brain have a range of membrane time constants due to varying properties such as dendritic morphology and the composition of conductances (see Section 3.1). As such, time constants can be individually optimized as parameters (similar to connectivity weights) [167, 168, 169, 170, 155, 154], or even parameterized to be input-dependent in gated network architectures (e.g., long short-term memory units, or LSTMs). Thus, by comparing to networks not endowed with diverse time constants, such studies provide evidence for the potential necessity of different timescales in computations.
For instance, several recent studies have shown that both recurrent and feedforward spiking networks trained to perform machine learning or cognitive tasks can benefit from both the heterogeneity and explicit optimization of single-neuron or spike-frequency adaptation time constants [114, 168, 84, 169, 171]. Strikingly, the trained networks exhibit similar timescale distributions as membrane time constants [168] or activity timescales of individual neurons in analogous brain regions of animals performing the same task [114]. Furthermore, time constants do not necessarily need to match that of the task, but both longer and diverse time constants were found to be beneficial [84]. Similarly, RNNs with trainable time constants demonstrate improved task performance and memory capacity while the learned time constants recover the hierarchy of timescales in the data-generating process [172, 173]. With more bio-realistic RNNs, such as those constrained by connectome measurements, training both the connection weights and single-neuron time constants resulted in emergent stimulus tuning similar to that of in vivo recorded neurons in the fly [170].
Remarkably, multi-layer feed-forward deep neural networks trained to simply mimic the input-output relationship of a single morphologically accurate pyramidal neuron recover complex and spatially varying temporal filters [174], further emphasizing the computational capacity provided by multiple timescales, even at the scale of single neurons. Recurrent architecture fitted to the same morphologically accurate pyramidal neuron allows a single neuron to solve tasks with complex long temporal dependencies [175]. On the other hand, diverse activity timescales can arise from network interactions instead of cellular properties (as discussed in the previous section). A recent study suggests that developing network-mediated diverse and long timescales may be a more robust mechanism for performing long memory tasks rather than trainable time constants [154]. Hence, future studies need to further distill the computational role of diverse cellular or network-mediated timescales.
These principles extend to more complex deep learning-based network architectures, providing functional benefits on more difficult tasks. Instead of trainable and fixed time constants, they can be further made to be input-dependent via trained gating weights (similar to gated RNNs, such as LSTMs and Gated Recurrent Units), which improves modeling of sequences with multiple timescales [176]. Furthermore, initializing gating parameters to reflect the distribution of timescales in the data improves performance on tasks involving memorizing and operating on long sequences [167]. Leveraging this observation, endowing LSTM units with a distribution of effective timescales constructed to match the power-law temporal dependency of natural language improves network performance in language modeling, especially long-range dependencies [177]. In parallel, these “chrono initialized” LSTMs improved decoding of brain response to narrative text, as well as estimation of voxel-wise fMRI BOLD timescales along the cortical hierarchy in human subjects [178], uniting potential principles of operation underlying both artificial and biological neural networks.
Finally, recent research in deep learning has shown an intriguing convergence of principles based on thoughtful inductive biases and initialization, as well as self-supervised pretraining, for improving performance on tasks with very long temporal dependencies [179]. Although RNN-based architectures can model infinitely long temporal dependencies in theory (similar to IIR filters), simpler implementations based on finite-length temporal convolutions appeared to be more effective at acquiring long timescale memories [180], though both architectures learn effective temporal context windows that reflect natural structures in data (e.g., speech [181]).
More recent architectures based on deep linear state space models (SSMs, e.g., LMU [182], S4 [183], and others [184, 185, 186, 187]) enjoy the benefits of both recurrence and convolutions, outperforming Transformer-based architectures on long timescale tasks, which are hindered by quadratic complexity in time. Remarkably, the success of these linear SSM-based models was in part driven by replacing the standard dense recurrent connections with structured and/or diagonal complex matrices, which directly parameterize the eigenvalues of the linear system—in other words, the system timescales (Box 3). Later extensions capitalize on these insights to derive better parameterization or initialization strategies, resulting in competitive models with simpler implementations, such as Structure Global Convolution [188] and Linear Recurrence Units [186]. On the other hand, it has been shown that both Transformer-based and SSM-based architectures can be drastically improved through self-supervised pretraining on the input sequence data, prior to downstream task training (e.g., labeled classification) on the same long-range tasks [189]. These findings suggest that pretraining or directly equipping deep learning models with inductive biases appropriate for long timescale tasks can be promising, but the exact mechanisms of these improvements—and to what extent they are related to intrinsic timescales in biological networks—remain unclear.
In summary, there is growing interest in convergent approaches to tuning network models to produce emergent task-relevant dynamics and timescales, as well as explicit optimization of time constant parameters in ANNs. Leveraging these approaches, findings from biological and deep learning network models suggest that both the magnitude and diversity of timescales may be important for tasks and data with long-range dependencies. Thus, these approaches appear to be a promising and collaborative avenue for research in neuroscience and AI aiming to understand and improve time-dependent computations in biological and artificial systems.
5 Discussion
We discussed how different computational methods help to turn empirical observation of diverse neural timescales into quantitative and testable hypotheses about brain dynamics: Different timescale estimation methods can disentangle various components of neural activity, such as modality- or task-specific variables, from intrinsic dynamics, while mechanistic models constrained by experimentally observed timescales, together with task-performing ANNs, can elucidate potential circuit mechanisms underlying brain dynamics and computations. Integrating these directions, we here provide an outlook for future research and discuss open questions.
5.1 Methodological suggestions
Given the differences in assumptions and applicability of the various timescale estimation methods we outlined, it is important to be aware of their limitations when interpreting consistencies and differences in existing results, as well as when designing future methodological and empirical research questions. Since there is not one single correct method, it would be ideal if each dataset was analyzed using all possible methods—a decidedly unrealistic proposal. Short of that, a systematic and quantitative comparison of different timescale estimation methods would serve as an extremely useful reference for contextualizing current and new results, while guiding the development of novel methods that address shared shortcomings of existing methods. Complementary to method development, neural timescale can be a standardized metric for characterizing neural dynamics at multiple scales, for which a curated database of existing estimates across brain regions, recording modalities, species, and task conditions—similar to databases of single-neuron electrophysiological measurements [190]—would be invaluable.
5.2 Models and data, an inseparable loop
Computational models suggest that various cellular and network mechanisms could alter timescales in the brain. Therefore, timescale-constrained models can be used as a great tool to provide mechanistic insights into brain dynamics and computations. It is important to note again that activity timescales of neurons may differ significantly from time constant parameters implemented in the models (e.g., membrane or synaptic time constants) (Box 2). Hence, it is crucial to match the activity timescales of the models to those of the data. For models with fewer mechanisms or free parameters, it is possible to derive the activity timescales analytically [104, 112]. However, in the brain, multiple mechanisms interact together to shape timescales, and studying their interactions analytically is not straightforward. In such cases, qualitative matching, e.g., generating a similar order of hierarchy of timescales [68, 106] or a grid search over multiple model parameters [17, 24] are used to fit the model timescales to data.
Nevertheless, such approaches may be inadequate when the parameter space is high-dimensional and the interaction between parameters is unknown. Fitting detailed mechanistic models with high-dimensional parameter space is possible using optimization or inference procedures such as genetic algorithms [191, 192], Approximate Bayesian Computation (ABC) [193, 194] and simulation-based inference (SBI) [26, 195, 196, 197, 198, 131, 27, 28]. In particular, since ABC and SBI estimate the posterior distribution of parameters, they identify degeneracies in parameter space and uncover potential relations between different parameters, which helps us better understand how combinations of mechanisms can give rise to observed neural timescales. Even with often imperfect (or, misspecified) models in neuroscience, when combined with inference methods that appropriately account for such errors [130], fitting mechanistic models to neural activity across multiple modalities [199, 200] could provide a holistic understanding of neural timescales across different scales of brain dynamics and may lead to new predictions that can be tested experimentally.
5.3 Timescales seed further synergies between neuroscience and AI
State-of-the-art machine learning tools have been readily adopted to interpret neural and behavioral data over multiple timescales. Conversely, the works discussed in Section 4 further attempt to build a theoretical bridge between principles of dynamic neural computation and deep and/or recurrent artificial neural networks. Here, we make two broad remarks that can guide future research.
First, viewing network timescale as the temporal signature of the geometry of its dynamical landscape, it stands to reason that optimizing the network to achieve a certain task-conducive geometry (e.g., line attractor or limit cycle) may be reformulated as optimization for a certain repertoire or diversity of timescales, as recent works have demonstrated a relationship between a network’s trainability to the stability of its dynamics (i.e., Lyapunov exponent) [162]. This has implications for RNN-based machine learning models, as “pretraining” for certain measures of network dynamic can be done without explicit task instructions [201, 202, 189, 155], and can also provide potential insights for how neural circuits organize themselves during early development without input [203, 204, 205].
Second, there is a growing interest in the emerging field of NeuroAI to develop models that not only can perform cognitive and behavioral tasks, but also produce realistic neural activity [206, 207, 208, 209]. While developing such models may seem challenging, matching models’ activity timescales to that of neural data can serve as a first step in building task-optimized models with realistic activity fluctuations. Timescale constraints may further alter the solution landscape of such models to what can be more relevant for understanding brain computations.
At the same time, modern AI models, such as large language models or deep reinforcement learning models, can facilitate studying the role of neural timescales in naturalistic tasks occurring over longer timescales than what can be studied in common laboratory-constrained experiments. Indeed, a recent study on freely behaving animals reported a different organization of neural timescales across brain areas than what has been found during resting state [41]. Therefore, timescales can serve as a key to open new avenues where ML tools can expedite uncovering mechanisms underlying brain computation.
5.4 Overlooked timescales and outlook
Much of systems and cognitive neuroscience has been based on a specific set of tasks, often with a trial structure between 1 to 10 seconds long, which artificially limits observed neural and behavioral timescale to a small range of the spectrum. For instance, spontaneous animal behavior in natural environments exhibits correlations over a long range of timescales that cannot be measured in common experimental settings [210, 211, 212]. On the other hand, processes such as the release of synaptic vesicles occur on a sub-second timescale [213] that are not detectable with many measurement tools. Furthermore, different physiological processes may be involved in shaping neural timescales that are often ignored when studying the mechanisms underlying neural timescales. For example, astrocytes have slow intracellular calcium fluctuations that may mediate slow neural modulations [214]. Astrocytes may also play a significant role in neural computations, as was recently shown that neuron–astrocyte networks can perform similar computations to complex machine learning models like Transformers [215]. How processes across different timescales interact may be of tremendous interest: does sub-second dynamics of synaptic vesicles play a role in shaping how, for example, memory engrams are formed over days and months, beyond the mean activity? On the other hand, do longer timescale fluctuations beyond a designated static state like age or disease conditions, such as circadian or seasonal rhythms, impact moment-to-moment neural dynamics? To investigate these questions, we likely need to leverage a combination of different experimental paradigms in the lab and naturalistic settings [210, 216, 217], together with long-term and high temporal resolution neural recording [218, 41], as well as novel data analysis and modeling tools to parse the interaction between timescales across different orders of magnitude.
Acknowlegments
This work was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC-Number 2064/1, PN 390727645) (RZ, JHM, RG) and SFB1233 (PN 276693517) (JHM), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039) (AL, JHM, RG), the Sofja Kovalevskaja Award from the Alexander von Humboldt Foundation endowed by the Federal Ministry of Education and Research (RZ, AL), the European Union (ERC, “DeepCoMechTome”, ref. 101089288) (JHM), the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101030918 (AutoMIND) (RG).
We thank Ana Manea and Jan Zimmermann for providing the timescales of fMRI data (Fig. 2a), Bradley Voytek and the speakers and participants of the Cosyne 2022 workshop on “Mechanisms, functions, and methods for diversity of neuronal and network timescales” for valuable discussions, and Scidraw (Federico Claudi and Andrew Hardaway), Wikipedia and Pixel perfect from Flaticon for providing the basis icons for some of the illustrations, and International Max Planck Research School for the Mechanisms of Mental Function and Dysfunction (IMPRS-MMFD) and Joachim Herz Foundation for their support.
Author contributions
RZ and RG conceptualized the overall structure of the review, wrote the first draft, and prepared the figures. All authors discussed, revised, and wrote the final version of the paper.